OpenStudy (lynn.caroline):

Please help with an explanation! For f(x) = 3x^2 - x - 2, find (f(x) - f(2))/(x-2)

OpenStudy (legomyego180):

Hello are you there?

OpenStudy (lynn.caroline):

Yes

OpenStudy (legomyego180):

\[f(x)=3x^2-x-2\] The best way to start this is by defining everything the equation is asking us to find. This is sort of a confusing question as well, its not asking for an answer, it just wants you to rewrite the equation.

OpenStudy (lynn.caroline):

So it doesn't want a real answer? How would I even got about rewriting the equation?

OpenStudy (legomyego180):

\[\frac{ f(x)-f(2) }{ x-2 }\] Lets define f(x)...the problem says: \[f(x)=3x^2-x-2\] So we can replace that into the function we are asked to find: \[\frac{ (3x^2-x-2)-f(2) }{ x-2 }\] Now we need to define f(2). To do this we just plug 2 in for everywhere there is an x in f(x): \[f(x)=3x^2-x-2\] \[f(2)=3(2)^2-2-2=8\] We can now plug in 8 for f(2): \[\frac{ (3x^2-x-2)-8 }{ x-2 }\] Simplify: \[\frac{ 3x^2-x-10 }{ x-2 }\]

OpenStudy (lynn.caroline):

So would the final answer be 3x+5?

OpenStudy (legomyego180):

Yup!

OpenStudy (legomyego180):

Good job. I forgot to factor the numerator.